# A Versatile Pathology Co-pilot via Reasoning Enhanced Multimodal Large Language Model

### v1

@@ 4.3 Curated Datasets
Rings. The Rings dataset [49] comprises WSIs of prostate biopsy specimens from 150 male patients, collected at the Division of Pathology, Department of Oncology (Turin, Italy). Tissue samples are formalin-fixed, paraffin-embedded, and sectioned, followed by hematoxylin and eosin (H&E) staining. Two expert pathologists manually annotated gland contours, categorizing each into two classes: prostate healthy glands and prostate tumor. The dataset includes 43,739 annotated ROIs, split into 29,451 training and 14,288 test samples.
SICAPv2. The SICAPv2 dataset [50] comprises 155 prostate biopsy WSIs obtained from 95 consenting patients, acquired using a Ventana iScan Coreo scanner at 40$\times$ magnification. Expert urogenital pathologists at Hospital Clínico of Valencia performed comprehensive histological analyses of the specimens. The dataset contains 23,924 annotated ROIs, strategically partitioned into 19,140 training samples and 4,784 test samples to ensure robust model development and evaluation.
TNBC. The TNBC dataset [51] comprises 167 annotated H&E-stained histopathology ROIs derived from triple-negative breast cancer specimens, systematically categorized into seven distinct cellular classes: glial, mitotic, cancerous, fibroblasts, lymphocyte plasmocyte, adipocytes, and endothelial. Captured at 40$\times$ magnification using a Philips Ultra Fast Scanner 1.6RA, the dataset spans eleven TNBC patients with 3-8 ROIs per case, deliberately sampling both tumor-rich regions and stromal microenvironments to ensure comprehensive representation of cellular diversity. Annotations are performed by a team of three experts (including one board-certified pathologist) through a rigorous multi-stage protocol: initial marking in ITK-SNAP with zoom-assisted nuclear boundary delineation, cross-validation by a second annotator, and final consensus resolution for disputed cases. The dataset is partitioned into 134 training ROIs and 33 test ROIs.
@@
PathMMU. The PathMMU dataset [55] represents a comprehensive multimodal resource for pathology VQA, aggregating data from five distinct sources: (1) PathMed (PubMed scientific articles), (2) Atlas (pathology textbooks), (3) EduContent (educational YouTube videos), (4) SocialPath (expert-contributed social media posts), and (5) PathCLS (existing pathology classification datasets). Due to access restrictions on SocialPath and training data, our study utilizes the official testtiny (555 samples) and validation (921 samples) subsets from the remaining four sources for evaluation, with 6,901 samples allocated for training. Detailed experimental results are provided in Extended Data Table 32.
@@
SlideInstrction-Caption. The SlideInstruction dataset [57] addresses the critical gap in large-scale multimodal resources for WSI understanding by providing 4,915 curated WSI-report pairs sourced from the TCGA database, spanning 4,028 patients. SlideInstrction-Caption Data, featuring concise and clinically focused summaries generated via GPT-4, highlights diagnostic findings while filtering noise (e.g., administrative text, specimen handling details).
@@
SlideInstrction-VQA. SlideInstrction-VQA comprises 175,753 VQA pairs structured into 3 broad classes (microscopy, diagnosis, clinical considerations) and 13 narrow subclasses (e.g., histopathological changes, differential diagnosis, and biomarker analysis ). The curation pipeline leverages GPT-4 to refine TCGA reports and generate diverse QA pairs while excluding low-relevance samples. To mimic real-world variability, the training set intentionally retains noisy cases where reports map ambiguously to multiple WSIs, enhancing model robustness. The test set strictly adheres to one-to-one WSI-report alignments for reliable evaluation. Detailed experimental results are provided in Extended Data Tables 40-42.
@@ 4.4 Compared Methods
Our study compares the performance of our model with five open-sourced multimodal models: Qwen2.5-VL, a general-purpose vision-language model; LLaVA-Med, a biomedical-specific model trained on a large corpus of medical literature; Quilt-LLaVA, a histopathology-focused model fine-tuned with spatially-grounded question-answer pairs; PathoR1, a contemporary work using reinforcement learning for reasoning refinement; and BiomedParse, a foundation model for joint segmentation and detection across various imaging modalities. By evaluating these models across a range of tasks, including ROI-level classification, detection, segmentation, VQA, WSI-level classification, and VQA, we aim to provide a comprehensive assessment of their strengths and weaknesses in computational pathology. This comparative analysis highlights the importance of domain-specific adaptations and the limitations of current methods in achieving robust generalizability across diverse histopathology applications.
Qwen2.5-VL. Qwen2.5-VL [25] constitutes a general-purpose vision-language model that advances multimodal understanding through several architectural innovations, including a dynamic resolution ViT encoder and optimized SwiGLU/RMSNorm components. The model demonstrates exceptional capability in document parsing and video temporal understanding, supported by its omnidocument processing framework that handles complex layouts, tables, and mathematical formulas. While not specifically optimized for medical applications, Qwen2.5-VL’s robust multilingual support and object grounding capabilities make it suitable for preliminary medical image analysis tasks, particularly in resource-constrained settings.
LLaVA-Med. LLaVA-Med [26] emerges as a biomedical multimodal large language model specifically designed for medical image understanding and reasoning, developed through curriculum learning on PubMed Central’s extensive collection of biomedical literature. The model architecture builds upon LLaVA’s foundation while incorporating domain-specific adaptations through two key phases: biomedical concept alignment using 1.6 million image-caption pairs from PMC-15M dataset, followed by instruction tuning with GPT-4 generated medical question-answer pairs. This approach enables LLaVA-Med to achieve state-of-the-art performance on biomedical VQA benchmarks.
@@ 4.5 Evaluation Metrics
where $N$ is the total number of samples,$\mathbb{I}()$ is the indicator function, $y_{i}$ and $\hat{y}_{i}$ denote the ground truth and predicted labels, respectively, and TP/TN/FP/FN represent true positives, true negatives, false positives, and false negatives.
@@ 4.6 Implementation Details
The model is implemented with PyTorch [65] and trained on a 8$\times$80GB H800 GPU node. $M$, the token count per pathology image, is set to 256 for ROI-level tasks and 1024 for WSI-level tasks, respectively. The patch size $P$ is set to 28. The intrinsic rank and global scaling factor in LoRA are set to 32 and 128. AdamW [66] is used as the optimizer with a weight decay of 0.1. The initial learning rate is set to 1e-4 for scale-dependent supervised fine-tuning and 1e-5 for task-aware reinforcement fine-tuning with a cosine learning rate schedule. KL regularization term coefficient $\beta$, clip coefficient $\epsilon$, and the number of sampled outputs $G$ in a group for GRPO are set to 0.001, 0.2, and 8, respectively.
@@
Z.X. and H.C. conceived and designed the work. Z.X., Z.L., J.H., J.M., C.J., Y.W., Z.C., Z.G., F.Z., Y.X., and X.W. curated the data included in the paper. Z.X. contributed to the technical implementation of the SmartPath-R1 framework and evaluated the performance of MLLMs. Z.X. wrote the manuscript with inputs from all authors. All authors reviewed and approved the final paper. H.C. supervised the research.

### v2

@@ 4.3 Curated Datasets
Rings. The Rings dataset [49] comprises WSIs of prostate biopsy specimens from 150 male patients, collected at the Division of Pathology, Department of Oncology (Turin, Italy). Tissue samples are formalin-fixed, paraffin-embedded, and sectioned, followed by hematoxylin and eosin (H&E) staining. Two expert pathologists manually annotated gland contours, categorizing each into two classes: prostate healthy glands and prostate tumors. The dataset includes 43,739 annotated ROIs, split into 29,451 training and 14,288 test samples.
[…]
@@
PathMMU. The PathMMU dataset [55] represents a comprehensive multimodal resource for pathology VQA, aggregating data from five distinct sources: (1) PathMed (PubMed scientific articles), (2) Atlas (pathology textbooks), (3) EduContent (educational YouTube videos), (4) SocialPath (expert-contributed social media posts), and (5) PathCLS (existing pathology classification datasets). Due to access restrictions on SocialPath and training data, our study utilizes the official testtiny and validation subsets from the remaining four sources for evaluation, with 6,901 samples allocated for training. Detailed experimental results are provided in Extended Data Table 32.
@@
SlideInstruction-Caption. The SlideInstruction dataset [57] addresses the critical gap in large-scale multimodal resources for WSI understanding by providing 4,915 curated WSI-report pairs sourced from the TCGA database, spanning 4,028 patients. SlideInstruction-Caption Data, featuring concise and clinically focused summaries generated via GPT-4, highlights diagnostic findings while filtering noise (e.g., administrative text, specimen handling details).
@@
SlideInstruction-VQA. SlideInstruction-VQA comprises 175,753 VQA pairs structured into 3 broad classes (microscopy, diagnosis, clinical considerations) and 13 narrow subclasses (e.g., histopathological changes, differential diagnosis, and biomarker analysis ). The curation pipeline leverages GPT-4 to refine TCGA reports and generate diverse QA pairs while excluding low-relevance samples. To mimic real-world variability, the training set intentionally retains noisy cases where reports map ambiguously to multiple WSIs, enhancing model robustness. The test set strictly adheres to one-to-one WSI-report alignments for reliable evaluation. Detailed experimental results are provided in Extended Data Tables 40-42.
@@ 4.4 Compared Methods
Our study compares the performance of our model with five open-sourced multimodal models: Qwen2.5-VL, a general purpose vision-language model; LLaVA-Med, a biomedical-specific model trained on a large corpus of medical literature; Quilt-LLaVA, a histopathology-focused model fine-tuned with spatially-grounded question-answer pairs; PathoR1, a contemporary work using reinforcement learning for reasoning refinement; and BiomedParse, a foundation model for joint segmentation and detection across various imaging modalities. By evaluating these models across a range of tasks, including ROI-level classification, detection, segmentation, VQA, WSI-level classification, and VQA, we aim to provide a comprehensive assessment of their strengths and weaknesses in computational pathology. This comparative analysis highlights the importance of domain-specific adaptations and the limitations of current methods in achieving robust generalizability across diverse histopathology applications.
Qwen2.5-VL. Qwen2.5-VL [25] constitutes a general purpose vision-language model that advances multimodal understanding through several architectural innovations, including a dynamic resolution ViT encoder and optimized SwiGLU/RMSNorm components. The model demonstrates exceptional capability in document parsing and video temporal understanding, supported by its omnidocument processing framework that handles complex layouts, tables, and mathematical formulas. While not specifically optimized for medical applications, Qwen2.5-VL’s robust multilingual support and object grounding capabilities make it suitable for preliminary medical image analysis tasks, particularly in resource-constrained settings.
LLaVA-Med. LLaVA-Med [26] emerges as a biomedical multimodal large language model specifically designed for medical image understanding and reasoning, developed through curriculum learning in the extensive collection of biomedical literature from PubMed Central. The model architecture builds upon LLaVA’s foundation while incorporating domain-specific adaptations through two key phases: biomedical concept alignment using 1.6 million image-caption pairs from PMC-15M dataset, followed by instruction tuning with GPT-4 generated medical question-answer pairs. This approach enables LLaVA-Med to achieve state-of-the-art performance on biomedical VQA benchmarks.
@@ 4.5 Evaluation Metrics
[…]
@@ 4.6 Implementation Details
The model is implemented with PyTorch [65] and trained on a $\times8$0GB H800 GPU node. $M$, the token count per pathology image, is set to 256 for ROI-level tasks and 1024 for WSI-level tasks, respectively. The patch size $P$ is set to 28. The intrinsic rank and global scaling factor in LoRA are set to 32 and 128. AdamW [66] is used as the optimizer with a weight decay of 0.1. The initial learning rate is set to 1e-4 for scale-dependent supervised fine-tuning and 1e-5 for task-aware reinforcement fine-tuning with a cosine learning rate schedule. KL regularization term coefficient $\beta$, clip coefficient $\epsilon$, and the number of sampled outputs $G$ in a group for GRPO are set to 0.001, 0.2, and 8, respectively.
@@
Z.X. and H.C. conceived and designed the work. Z.X., Z.L., J.H., J.M., C.J., Y.W., Z.C., Z.G., F.Z., and Y.X., curated the data included in the paper. Z.X. contributed to the technical implementation of the SmartPath-R1 framework and evaluated the performance of MLLMs. Z.X. wrote the manuscript with inputs from all authors. All authors reviewed and approved the final paper. H.C. supervised the research.